OpenStudy (anonymous):

A principal of $3900 is invested at 3.5% interest, compounded annually. How much will the investment be worth after 7 years? Use the calculator provided and round your answer to the nearest dollar.

OpenStudy (gorv):

\[C.I=P*(1+\frac{ r }{ 100 })^t\]

OpenStudy (gorv):

r=3.5 t=7

OpenStudy (gorv):

after 7 year investment= 3900 + C.I

OpenStudy (anonymous):

After putting that last piece together I got 1.00492659.

OpenStudy (anonymous):

THE ANSWER IS $4962 =·39001.03574961.88… Rounded to the nearest dollar, we get $4962 .

OpenStudy (gorv):

that is your compound interest sum=3900+C.I